OpenStudy (anonymous):

if \[\sqrt{4-\sqrt{8+\sqrt{32{+\sqrt{768}}}}}=a \sqrt{2}\cos \frac{ 11 \pi }{ b }\] find a and b!

OpenStudy (anonymous):

@abb0t ? @preteenpony ? @ganeshie8 ? @hugsnotughs ?

OpenStudy (anonymous):

@ganeshie8 ?@Abhisar ? @bradely ? @Compassionate ? @D3xt3R ?

OpenStudy (anonymous):

@Owlcoffee ? @quickstudent ? @radar ? @SolomonZelman ?

OpenStudy (anonymous):

Here's something to notice: \[\begin{align*}\sqrt{4-\sqrt{8+\sqrt{32{+\sqrt{768}}}}}&=a \sqrt{2}\cos \frac{ 11 \pi }{ b }\\ \sqrt{\frac{1}{2}}\sqrt{4-\sqrt{8+\sqrt{32{+\sqrt{768}}}}}&=a\cos \frac{ 11 \pi }{ b }\\ \sqrt{2-\frac{1}{2}\sqrt{8+\sqrt{32{+\sqrt{768}}}}}&=a\cos \frac{ 11 \pi }{ b }\\ \sqrt{2-\sqrt{\frac{8}{4}+\frac{1}{4}\sqrt{32{+\sqrt{768}}}}}&=a\cos \frac{ 11 \pi }{ b }\\ \sqrt{2-\sqrt{2+\sqrt{\frac{32}{16}+\frac{1}{16}\sqrt{768}}}}&=a\cos \frac{ 11 \pi }{ b }\\ \sqrt{2-\sqrt{2+\sqrt{2+\sqrt{\frac{768}{256}}}}}&=a\cos \frac{ 11 \pi }{ b }\\ \sqrt{2-\sqrt{2+\sqrt{2+\sqrt3}}}&=a\cos \frac{ 11 \pi }{ b } \end{align*}\] Not sure if that helps though...

OpenStudy (quickstudent):

Sorry, I don't know how to do this.

OpenStudy (anonymous):

@mathstudent55 ?

OpenStudy (anonymous):

@ganeshie8 ?

OpenStudy (anonymous):

What if you use the fact that \(-1\le\cos\dfrac{11\pi}{b}\le1\)? You might be able to find some bounds for \(a\).

OpenStudy (anonymous):

yes i guess so what should be our approach now?

OpenStudy (anonymous):

Oh I have no idea, just throwing out some ideas ;)

ganeshie8 (ganeshie8):

i am trying to do the impossible task of simplifying left hand side \(\large \sqrt{2+\sqrt{3}} = \dfrac{\sqrt{2} + \sqrt{6}}{2}\) next level gives me 3 terms and i feel stuck

ganeshie8 (ganeshie8):

\[\large \sqrt{2-\frac{1}{4}\sqrt{4+\sqrt{2} + \sqrt{6} }} =a\cos \frac{ 11 \pi }{ b } \]

ganeshie8 (ganeshie8):

@mukushla

OpenStudy (anonymous):

Is it even possible to find values for two unknowns with just the one equation? Or doe the periodic part have something to do with it?

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

@ganeshie8 ?

OpenStudy (anonymous):

@abb0t ? @Compassionate ? @dan815 ? @esshotwired ?

OpenStudy (anonymous):

@dan815 ?

OpenStudy (xapproachesinfinity):

can you use the same method that @ganeshie8 used repeatedly for the rest of the root I dunno what that will give you